OpenStudy (anonymous):

Solve the system, using substitution. What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n A. 2 B. 0 C. -4 D. 2 E. no solution

OpenStudy (anonymous):

2(3-2n) - 3n = 6 6 - 4n -3n = 6 6 - n = 6 n = 12? Check to see if I did it right

OpenStudy (a1234):

You subtract 6, not add.

OpenStudy (anonymous):

Right! Thanks. n = 6 - 6 = 0. B is your answer.